The aspect ratio of a TV screen is the ratio of its length to its width. The aspect ratio of a standard TV screen is 4:3. The aspect ratio of a wide screen TV in United States is 16:9. Describe how you would tell whether a TV has a standard or wide screen given the length and width of the screen.
Answer by checkley71(8403)

IF 4/3=16/9 THEN THE LARGER TV IS STANDARD
WE CROSS MULTIPLY
3*16=4*9
48 DOES NOT =36. THEREFORE THE LARGER SCREEN IS NOT STANDARD.